OpenStudy (anonymous):

Find the area of a regular hexagon with the given measurement. 6-inch radius A = sq. in.

OpenStudy (anonymous):

i want to help you get to the answer on your own, do you know an equation for the area of a regular polygon?

OpenStudy (anonymous):

I just started 10.1 and a little lost

OpenStudy (anonymous):

ok, so online i found a general equation for regular polygons (which are hexagons, pentagons etc). it is: \[\frac{ 1 }{ 2 }*n*r^2*\sin (\frac{ 360 }{ n })\] where r is radius and n is the number of sides you have. a hexagon has 6 sides, and you have a 6inch radius. so based on that information the formula is just plug and chug

OpenStudy (anonymous):

Okay thank you. I will give it a try

OpenStudy (anonymous):

see if you can find the answer and let me know so i can compare it to the answer i got

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

Hello???? Still confused

OpenStudy (anonymous):

im sorry im sorry, i was working on my own things. go ahead and put that equation you had back up there because i believe it was correct and it can be a good ground work to start from

OpenStudy (anonymous):

1/2*n*6^2*sin (360/n)

OpenStudy (anonymous):

ok then what

OpenStudy (anonymous):

n is the number of sides in the shape, what is n for this problem?

OpenStudy (anonymous):

I don't know

OpenStudy (anonymous):

there are 6 sides in a hexagon

OpenStudy (anonymous):

I know that

OpenStudy (anonymous):

right, so help me understand what you dont understand then? because all you have to do is plug 6in for R and 6 into n in the equation and solve it

OpenStudy (anonymous):

You are given the Hexgaon below.Given that the radius is 6 inches, find the area of the Hexagon. Explain in detail how you found the area. HINT: Can you find the area of a triangle? How many triangles would there be? (This is what my extra credit says)

OpenStudy (anonymous):

finding the triangles is just another way to do it. and you did not specify that is the method you wanted to use, you only said you needed area so that is what i solved for. but there would be 6 triangles

OpenStudy (anonymous):

Okay let me plug in 6 for both R and N

OpenStudy (anonymous):

let me know what you get

OpenStudy (anonymous):

I got 93.5

OpenStudy (anonymous):

that is the same thing i got. so your answer is correct

OpenStudy (anonymous):

Okay. Thank you

OpenStudy (anonymous):

no problem